# Detecting and Quantifying Isolated Singularities over Discrete Valuation Rings

### v1

@@ Abstract
This paper develops a theory of isolated hypersurface singularities in mixed characteristic $(0,p)$, focusing on quotient rings over a Discrete Valuation Ring (DVR). We introduce and study analogues of the classical Tjurina and Milnor numbers for this setting, prove a generalized analogue of the Determinacy Theorem and the Mather-Yau Theorem for complete Noetherian local rings, and define numerical invariants that provide distinct criteria for detecting isolated singularities in the unramified and ramified cases.
@@ 1 Introduction
The studies of isolated singularities originated with the foundational work of Arnold in [varchenko1985singularities], Milnor in [milnor2016singular], and Tjurina in [tjurina1969locally] in the complex setting. Their pioneering efforts involved the computation and analysis of hypersurface singularities using numerical invariants, specifically the Milnor and Tjurina numbers, along with key concepts like equivalences and determinacy. This framework was later generalized by Greuel and Kröning in [greuel1990simple] to algebraically closed fields of arbitrary characteristic, with subsequent expansions by Boubakri in [boubakri2009hypersurface] and Nguyen in [Nguyen2013classification].
This paper continues this line of generalization by extending the theory of isolated singularities of hypersurfaces from algebraically closed fields to a more general setting of hypersurfaces over the ring of power series over a local ring. Specifically, we focus on quantifying and understanding quotient rings $\frac{V[[\underline{x}]]}{\langle f\rangle}$, where $V$ is a Discrete Valuation Ring (DVR) with a chosen uniformizer $\pi$, with isolated singularities, i.e. such that $\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)_{\mathfrak{q}}$ is regular for every non-maximal prime ideal $\mathfrak{q}$. The key challenge lies in developing numerical invariants, analogous to the classical Tjurina and Milnor numbers, that can effectively detect and measure the severity of these singularities.
When moving to the DVR case, there is a constraint that we do not encounter when working over fields. In this case, the uniformizer $\pi$ of $V$ is non invertible, and behaves differently than the units in $V$. This causes the groups of $V$-automorphisms of the ring of formal power series over $V$ to be much smaller compared to its field counterpart. Therefore elements $f,g\in V[[\underline{x}]]$ that we would want to be considered as "equivalent up to a change of variables" can not be. For example, as in Example 1.1. in [carvajal2019covers], one would want to view $x_{1}^{2}+x_{2}^{2}+\pi^{3}$ and $x_{1}^{3}+x_{2}^{2}+\pi^{2}$ as equivalent. But we have that
@@
This point of view allows us to formally define an analogue of the Tjurina and Milnor numbers over $V$ that behave similarly their field counterparts. In addition, it allows us to define equivalences $\sim$ of hypersurface singularities (defined formally in Definition 3.9). Using the notion of equivalence and the Tjurina number over $V$, we develop a theory of hypersurface singularities over $V$, analogous to the results over fields, as summarized in Section 2 of [greuel2007introduction]. Our main results include the following:
@@
    We present a generalization of the Determinacy Theorem (Section 2) to complete Noetherian local rings, inspired by Pellikaan [pellikaan1988finite] and Boubakri-Greuel-Markwig in [boubakri2012invariants]. This result is then used to prove a generalized Mather-Yau theorem and to define a new numerical invariant for singularities over complete local rings that measures the determinacy.
@@
    We define numerical invariants (in Section 4) to detect isolated singularities based on whether $V$ is ramified or unramified, drawing on recent works by Hochster and Jefferies [hochster2021jacobian], Saito [saito2022frobenius], and KC [kc2024singular].
We summarize these results in the following Theorem, which is a combination of Theorem 2.4, Proposition 2.28, Proposition 3.14, Proposition 3.27, Proposition 4.14, Corollary 4.23, and Proposition 4.27:
Theorem.
@@
      Assume that $f$ has finite determinacy with respect to $I$ (as in Definition 2.1). Then $f\overset{I}{\sim}g$ (as in Definition 2.1) if and only if for every $k\gg 0$ we have an isomorphism of $R$-algebras
@@
      $\tau_{V}(f)=0$ (as in Definition 3.4) if and only if $f\sim x_{1}$ if and only if $\frac{V[[\underline{x}]]}{\langle f\rangle}$ is a regular local ring.
@@
      If the characteristic of $\kappa$ is not 22 and $\kappa$ is quadratically closed then $\tau_{V}(f)=1$ if and only if $f\sim\pi^{2}+x_{1}^{2}+\cdots+x_{n}^{2}$.
@@
      If $V$ is unramified then $f$ defines an isolated singularity if and only if $p\in\sqrt{J(f)}$ and $\tau(f,\delta)<\infty$ for every $p$-derivation $\delta$ (as in Definition 4.11) if and only if $p\in\sqrt{J(f)}$ and $\tau^{\Delta}(f)<\infty$ (as in Corollary 4.23).
@@
      If $V$ is ramified then $f$ defines an isolated singularity if and only if $\pi\in\sqrt{J(f)}$ and $\tau^{\pi}(f)<\infty$ (as in Definition 4.26).
@@
The second, $\tau(f,\delta)$, is defined in Definition 4.11 over unramified DVRs with respect to a chosen $p$-derivation $\delta$, which we can view as a operator that "derives with respect to pp". $\tau(f,\delta)$ detects isolated singularities completely (with an additional technical condition), yet it behaves very differently than the Tjurina number as defined over fields, in addition to highly depending on the choice of $\delta$. The third, $\tau^{\Delta}(f)$, is the value of $\tau(f,\delta)$ when going $\operatorname{mod}\,p$, does not depend on the choice of $\delta$ and behaves similarly to $\tau(f,\delta)$.
@@
Notations: Given a local ring $(R,\mathfrak{m},\kappa)$, we denote by $R[[\underline{x}]]=R[[x_{1},\dots,x_{n}]]$ the ring of power series over $R$ in variables $x_{1},\dots,x_{n}$ and we denote its maximal ideal by $\mathfrak{M}=\mathfrak{m}+\langle x_{1},\dots,x_{n}\rangle$. Note that $R[[\underline{x}]]$ is a complete topological ring with respect to the $\mathfrak{M}$-adic topology. Given $f,g\in R[[\underline{x}]]$, we denote $f\overset{\mathcal{K}}{\sim}g$ if we have an isomorphism of $R$-algebras $\frac{R[[\underline{x}]]}{\langle f\rangle}\cong\frac{R[[\underline{x}]]}{\langle g\rangle}$, and we denote $f\overset{\mathcal{R}}{\sim}g$ if there exists some $R$-automorphism $\varphi$ such that $\varphi(f)=g$. Given some $f$ and some variable $z$, we denote by $\partial_{z}(f)$ the partial derivative of $f$ with respect to $z$, we denote $\partial_{i}(f)=\partial_{x_{i}}(f)$, and we denote $J(f)=\langle\partial_{1}(f),\dots,\partial_{n}(f)\rangle$. Given an ideal $I$, we denote by $\operatorname{ord}_{I}(f)$ the largest $k\geq 0$ such that $f\in I^{k}$ and we denote $\operatorname{ord}(f)=\operatorname{ord}_{\mathfrak{M}}(f)$. In addition, given some module $M$ over $R$, we denote by $\textup{length}_{R}\left(M\right)$ the length of the module $M$ over $R$. When discussing power series, we use multi-index notation (as in [greuel2007introduction]): If $\alpha=(\alpha_{1},\dots,\alpha_{n})$ is a tuple of non negative integers and $\underline{x}=(x_{1},\dots,x_{n})$ is a tuple of ring elements, we denote $|\alpha|=\alpha_{1}+\cdots+\alpha_{n}$, we denote $\underline{x}^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$, and we denote $\sum_{\alpha}a_{\alpha}\underline{x}^{\alpha}=\sum_{\alpha_{1},\dots,\alpha_{n}\geq 0}a_{\alpha_{1},\dots,\alpha_{n}}x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$. Given $\alpha=(\alpha_{1},\dots,\alpha_{n})$ and $\beta=(\beta_{1},\dots,\beta_{n})$, we say that $\beta\geq\alpha$ if $\beta_{j}\geq\alpha_{j}$ for every $j=1,\dots,n$. Given $f(\underline{x})=\sum_{\alpha}a_{\alpha}\underline{x}^{\alpha}$, we denote $f(\underline{0})=a_{0}$.
Acknowledgments. This work was done as part of the Author’s PhD thesis under the guidance of Karl Schwede, and we wish to thank him for his guidance, help, and support. We wish to thank Neil Epstein, Gert-Martin Greuel, Sean Howe, Jack Jeffries, Nawaj KC, Hong Duc Nguyen, and Eugenii Shustin for productive mathematical discussions and their inputs on some of the ideas presented in this paper. The author was partially support by NSF grant DMS-2101800.
@@ 2 Determinacy over Complete Local Rings
In this section we provide a proof of the finite determinacy theorem over any complete Noetherian local ring $(R,\mathfrak{m},\kappa)$. These determinacy results are generalized versions of the results presented in Section 3.1 of [boubakri2009hypersurface] (which in turn are based on Section 2.2 of [greuel2007introduction] in the characteristic zero case and were expanded upon in Section 2 of [boubakri2012invariants] and in [greuel2019finite]).
@@
Assume that $R$ is a DVR with uniformizer $\pi$, that $\kappa$ is quadratically closed, and set $f=\pi^{2}+x^{2}\in R[[x]]$. We show directly that $f$ is $2$-determined with respect to $\mathfrak{M}=\langle x,\pi\rangle$. Given some $g\in R[[x]]$ such that $f-g\in\mathfrak{M}^{3}$, we can write $g=x^{2}+\pi^{2}+h$ where $h\in\langle x,\pi\rangle^{3}$, and so $h=\alpha_{1}x^{3}+\alpha_{2}x^{2}\pi+\alpha_{3}x\pi^{2}+\alpha_{4}\pi^{3}$ for some $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}\in R[[\underline{x}]]$. Therefore we have that
@@
Observe that both $1+\alpha_{1}x+\alpha_{2}\pi$ and $1+\alpha_{3}x+\alpha_{4}\pi$ are units. Therefore, since $R$ is a DVR (and therefore Henselian) and since $\kappa$ is quadratically closed, there exists some unit $\beta\in R[[\underline{x}]]$ such that $\beta^{2}=(1+\alpha_{1}x+\alpha_{2}\pi)(1+\alpha_{3}x+\alpha_{4}\pi)^{-1}$. So, if we define the automorphism $\varphi$ by setting $\varphi(x)=\beta x$ then we can conclude that $\varphi(f)=(1+\varphi(\alpha_{3})\beta x+\varphi(\alpha_{4})\pi)g$ with $1+\varphi(\alpha_{3})\beta x+\varphi(\alpha_{4})\pi$ being a unit, and thus $f\overset{\mathcal{K}}{\sim}g$, as desired.
@@
Determinacy over local rings that are algebras over an algebraically closed field of characteristic zero have been studied by Belitskii and Kerner in [belitskii2012matrices, belitskii2016finite, belitskii2019finite]. For more information on determinacy of general group actions, see [boix2022pairs]. In addition, the study of the group $\mathcal{D}_{I}$ where $I$ is an ideal in the ring of power series over $\mathbb{C}$ was presented in [pellikaan1988finite].
@@
  • The $R$-homomorphism $\varphi$ defined by $\varphi(x_{j})=x_{j}+b_{j}$ for every $j$ is an $R$-automorphism of $R[[\underline{x}]]$.
@@
Note that we can rewrite Lemma 2.9 as $\operatorname{ord}_{I}(\partial_{i}(f))\geq\operatorname{ord}_{I}(f)-1$ for every ideal $I\subset R[[\underline{x}]]$. Note that if $R$ is a ring of either mixed or positive characteristic then the order of the derivative can substantially change and is unbounded from above. For example, if $R$ is of mixed characteristic $(0,p)$ and $f(x)=x^{p^{r}}\in R[[x]]$ for some $r>1$ then we have that $\text{ord}(\partial_{x}(f))=\text{ord}(f)+r-1$.
@@
    (Implicit Function Theorem): Given some $f_{1},\dots,f_{m}\in R[[\underline{x},y_{1},\dots y_{m}]]$ such that $f_{i}(\underline{0})=0$ for every $i$ and
@@
    then there exists unique $g_{1},\dots,g_{m}\in\mathfrak{M}\subset R[[\underline{x}]]$ for which we have that $f_{i}\left(x_{1},\dots,x_{n},g_{1},\dots g_{m}\right)=0$ for every $i$.
@@
    (Inverse Function Theorem): Let $f_{1},\dots,f_{n}\in\mathfrak{M}\subset R[[\underline{x}]]$. Then the Jacobian matrix $Df=\left[\frac{\partial f_{i}}{\partial x_{j}}\left(0\right)\right]_{i,j}\in\operatorname{Mat}_{n}\left(R\right)$ is invertible if and only if the unique $R$-homomorphism $\varphi\colon R[[\underline{x}]]\to R[[\underline{x}]]$ defined by $\varphi\left(x_{i}\right)=f_{i}$ is an $R$-automorphism.
@@
    Since $R[[\underline{x}]]$ is a complete local ring, then by [stacks-project, Tag 04GM] (or by corollary 1.9 in [leuschke2012cohen]) it must be Henselian. Therefore, given $f_{1},\dots,f_{m}\in R[[\underline{x},y_{1},\dots,y_{m}]]=R[[\underline{x}]][[y_{1},\dots,y_{m}]]$ such that satisfy the condition above, then by setting $\overline{f_{i}}=f_{i}\,\operatorname{mod}\,\mathfrak{M}\in\kappa[[y_{1},\dots,y_{m}]]$ for every $i$ we must have that $\overline{f_{i}}\in\langle y_{1},\dots,y_{m}\rangle\subset\kappa[[\underline{y}]]$ and that the Jacobian matrix of $(\overline{f_{1}},\dots,\overline{f_{m}})$ is an invertible matrix over $\kappa$. Therefore, by the Implicit function theorem over $\kappa$ and by the Multi-dimensional Hensel’s lemma (see, for example, Section 4 in [kuhlmann2000valuation], Section 7 in [ziegler1978model], or Section 1.2 of Chapter I in [greuel2007introduction]), we can find some unique $\overline{g_{1}},\dots,\overline{g_{m}}$ such that $\overline{f_{i}}\left(x_{1},\dots,x_{n},\overline{g_{1}},\dots,\overline{g_{m}}\right)=0$ for every $i$ that uniquely lift to some $g_{1},\dots,g_{m}\in R[[\underline{x},y_{1},\dots y_{m}]]$ such that $f_{i}\left(x_{1},\dots,x_{n},g_{1},\dots,g_{m}\right)=0$ for every $i$, as desired.
@@
In addition, if $\varphi$ is an $R$-isomorphism then we can construct $\Phi$ to be an $R$-isomorphism as well with $\varphi(I)=J$.
@@
Note that Lemma 2.13 and Lemma 2.14 give us analogues to some classical results in the field case. The first item of Lemma 2.13 is a generalized version of [tougeron1968ideaux] (for more details, see Section 3.1 in [hauser2013artin], Section 3 in [belitskii2016strong], or [kerner2024orbits]) and the second item is a version of Corollary 3.3.7 in [de2013local]. Lemma 2.14 is a generalized version of Lemma 1.14 and Lemma 1.23 in [greuel2007introduction] and of Lemma 1.3.5. in [boubakri2009hypersurface].
@@
then we have that $F(\underline{x},\varphi_{i}(\underline{x}),u_{i})=f-u_{i}g\circ\varphi_{i}\to 0$. So by the Artin’s strong approximation theorem over local rings (see, for example, Theorem 3.3. in [rond2015artin], which in turn cites [schappacher1980some, schappacher1983inegalite]) we have that there exists some automorphism $\varphi$ and some unit $u$ such that $u_{i}\varphi_{i}(\underline{x})-u\varphi(\underline{x})\in\mathfrak{M}^{i}$ and $F(\underline{x},\varphi(\underline{x}),u)=0$, which gives us that $f=ug(\varphi)$. ∎
@@
Artin’s approximation theorem that we use in the proof of Lemma 2.16, which is based upon Artin’s original result in [artin1969algebraic], is an essential result in the theory of isolated singularities over fields. For more information and details, see [belitskii2020approximation, hauser2013artin].
@@
The proof of Lemma 2.18 is similar to that of Lemma 1.2.7. in [boubakri2009hypersurface] and to that of Lemma 2.10. in [greuel2007introduction].
@@
which must be contained in $\mathfrak{M}^{N+2}$. Now, we can proceed inductively (using the notations of Lemma 2.7) to get collection of sequences $\{\{b_{i,j}\}_{i=1}^{\infty}\colon j=1,\dots,n\}$ such that $b_{i,0}\in\mathfrak{M}^{N+i-1}$ and $b_{i,j}\in\mathfrak{M}^{N+i}$ for every $j=1,\dots,n$. Therefore by setting $\varphi_{i}(x_{j})=x_{j}+b_{i,j}$ for every $i,j$, we get that $g-v_{i}\cdot\varphi_{i}(f)\in\mathfrak{M}^{N+1+i}$ and that $\varphi_{i}\in\mathcal{D}_{I}$, as desired. ∎

### v2

@@ Abstract
We develop a theory of isolated hypersurface singularities in mixed characteristic $(0,p)$, focusing on quotient rings over a Discrete Valuation Ring (DVR). We introduce and study analogues of the classical Tjurina and Milnor numbers for this setting, prove a generalized analogue of the Determinacy Theorem and the Mather-Yau Theorem for complete Noetherian local rings, and define numerical invariants that provide distinct criteria for detecting isolated singularities in the unramified and ramified cases.
@@ 1 Introduction
The studies of hypersurfaces with isolated singularities originated with the foundational work of Arnold in [VG85], Milnor in [MIL16], and Tjurina in [TJU69] in the complex setting. Their pioneering efforts involved the computation and analysis of hypersurface singularities using numerical invariants, specifically the Milnor and Tjurina numbers, along with key concepts like equivalences and determinacy. This framework was later generalized by Greuel and Kröning in [GK90] to algebraically closed fields of arbitrary characteristic, with subsequent expansions by Boubakri in [BOU09] and Nguyen in [NGU13].
This chapter continues this line of generalization by extending the theory of isolated singularities of hypersurfaces from algebraically closed fields to a more general setting of hypersurfaces over the ring of power series over a local ring. Specifically, we focus on quantifying and understanding quotient rings $\frac{V[[\underline{x}]]}{\langle f\rangle}$, where $V$ is a Discrete Valuation Ring (DVR) with a chosen uniformizer $\pi$, with isolated singularities, i.e. such that $\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)_{\mathfrak{q}}$ is regular for every non-maximal prime ideal $\mathfrak{q}$. The key challenge lies in developing numerical invariants, analogous to the classical Tjurina and Milnor numbers, that can effectively detect and measure the severity of these singularities.
When moving to the DVR case, there is a constraint that we do not encounter when working over fields. In this case, the uniformizer $\pi$ of $V$ is non invertible, and behaves differently than the units in $V$. This causes the groups of $V$-automorphisms of the ring of formal power series over $V$ to be much smaller compared to its field counterpart. Therefore elements $f,g\in V[[\underline{x}]]$ that we would want to be considered as "equivalent up to a change of variables" can not be. For example, as in Example 1.1. in [CMP+21], one would want to view $x_{1}^{2}+x_{2}^{2}+\pi^{3}$ and $x_{1}^{3}+x_{2}^{2}+\pi^{2}$ as equivalent. But we have that
@@
This point of view allows us to formally define an analogue of the Tjurina and Milnor numbers over $V$ that behave similarly their field counterparts. In addition, it allows us to define equivalences $\sim$ of hypersurface singularities (defined formally in Definition 3.9). Using the notion of equivalence and the Tjurina number over $V$, we develop a theory of hypersurface singularities over $V$, analogous to the results over fields, as summarized in Section 2 of [GLS07]. Our main results include the following:
@@
    We present a generalization of the Determinacy Theorem (Section 2) to complete Noetherian local rings, inspired by Pellikaan [PEL88] and Boubakri-Greuel-Markwig in [BGM12]. This result is then used to prove a generalized Mather-Yau theorem and to define a new numerical invariant for singularities over complete local rings that measures the determinacy.
@@
    We define numerical invariants (in Section 4) to detect isolated singularities based on whether $V$ is ramified or unramified, drawing on recent works by Hochster and Jeffries in [HJ21], Saito in [SAI22], and KC in [KC24].
We summarize these results in the following Theorem:
Theorem 1.1.
@@
      (Proposition 2.26) If $f$ has finite determinacy with respect to $I$ then $\textup{Sing}(V(f))\subset V(I+\langle f\rangle+J(f))$.
    3. (c)
      (Proposition 2.33) Assume that $f$ has finite determinacy with respect to $I$. Then $f\overset{I}{\sim}g$ (as in Definition 2.1) if and only if for every $k\gg 0$ we have an isomorphism of $R$-algebras
@@
      (Proposition 3.14) $\tau_{V}(f)=0$ (as in Definition 3.4) if and only if $f\sim x_{1}$ if and only if $\frac{V[[\underline{x}]]}{\langle f\rangle}$ is a regular local ring.
@@
      (Proposition 3.27) If the characteristic of $\kappa$ is not $2$ and $\kappa$ is quadratically closed then $\tau_{V}(f)=1$ if and only if $f\sim\pi^{2}+x_{1}^{2}+\cdots+x_{n}^{2}$.
@@
      (Proposition 4.14, Corollary 4.22) If $V$ is unramified then $f$ defines an isolated singularity if and only if $p\in\sqrt{J(f)}$ and $\tau(f,\delta)<\infty$ for every $p$-derivation $\delta$ (as in Definition 4.11) if and only if $p\in\sqrt{J(f)}$ and $\tau^{\Delta}(f)<\infty$ (as in Corollary 4.22).
@@
      (Proposition 4.26) If $V$ is ramified then $f$ defines an isolated singularity if and only if $\pi\in\sqrt{J(f)}$ and $\tau^{\pi}(f)<\infty$ (as in Definition 4.25).
@@
[…]
@@
Notations: Given a local ring $(R,\mathfrak{m},\kappa)$, we denote by $R[[\underline{x}]]=R[[x_{1},\dots,x_{n}]]$ the ring of power series over $R$ in variables $x_{1},\dots,x_{n}$ and we denote its maximal ideal by $\mathfrak{M}=\mathfrak{m}+\langle x_{1},\dots,x_{n}\rangle$. Note that $R[[\underline{x}]]$ is a complete topological ring with respect to the $\mathfrak{M}$-adic topology. Given an ideal $\mathfrak{a}\subset R[[\underline{x}]]$, we denote by $V(\mathfrak{a})$ the set of all prime ideals $\mathfrak{a}\subset\mathfrak{p}$. Given $f,g\in R[[\underline{x}]]$, we denote $f\overset{\mathcal{K}}{\sim}g$ if we have an isomorphism of $R$-algebras $\frac{R[[\underline{x}]]}{\langle f\rangle}\cong\frac{R[[\underline{x}]]}{\langle g\rangle}$, and we denote $f\overset{\mathcal{R}}{\sim}g$ if there exists some $R$-automorphism $\varphi$ such that $\varphi(f)=g$. Given some $f$ and some variable $z$, we denote by $\partial_{z}(f)$ the partial derivative of $f$ with respect to $z$, we denote $\partial_{i}(f)=\partial_{x_{i}}(f)$, and we denote $J(f)=\langle\partial_{1}(f),\dots,\partial_{n}(f)\rangle$. Given an ideal $I$, we denote by $\operatorname{ord}_{I}(f)$ the largest $k\geq 0$ such that $f\in I^{k}$ and we denote $\operatorname{ord}(f)=\operatorname{ord}_{\mathfrak{M}}(f)$. In addition, given some module $M$ over $R$, we denote by $\textup{length}_{R}\left(M\right)$ the length of the module $M$ over $R$. When discussing power series, we use multi-index notation (as in [GLS07]): If $\alpha=(\alpha_{1},\dots,\alpha_{n})$ is a tuple of non negative integers and $\underline{x}=(x_{1},\dots,x_{n})$ is a tuple of ring elements, we denote $|\alpha|=\alpha_{1}+\cdots+\alpha_{n}$, we denote $\underline{x}^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$, and we denote $\sum_{\alpha}a_{\alpha}\underline{x}^{\alpha}=\sum_{\alpha_{1},\dots,\alpha_{n}\geq 0}a_{\alpha_{1},\dots,\alpha_{n}}x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$. Given $\alpha=(\alpha_{1},\dots,\alpha_{n})$ and $\beta=(\beta_{1},\dots,\beta_{n})$, we say that $\beta\geq\alpha$ if $\beta_{j}\geq\alpha_{j}$ for every $j=1,\dots,n$. Given $f(\underline{x})=\sum_{\alpha}a_{\alpha}\underline{x}^{\alpha}$, we denote $f(\underline{0})=a_{0}$.
Acknowledgments. This work was done as part of the Author’s PhD thesis under the guidance of Karl Schwede, and we wish to thank him for his guidance, help, and support. We wish to thank Neil Epstein, Gert-Martin Greuel, Sean Howe, Jack Jeffries, Nawaj KC, Hong Duc Nguyen, and Eugenii Shustin for productive mathematical discussions and their inputs on the different ideas presented. We especially wish to thank Benjamin Baily for the main idea and for the technical details of Proposition 2.26. The author was partially support by NSF grant DMS-2101800.
@@ 2 Determinacy over Complete Local Rings
In this section we provide a proof of the finite determinacy theorem over any complete Noetherian local ring $(R,\mathfrak{m},\kappa)$. These determinacy results are generalized versions of the results presented in Section 3.1 of [BOU09] (which in turn are based on Section 2.2 of [GLS07] in the characteristic zero case and were expanded upon in Section 2 of [BGM12] and in [GP19]).
@@
Assume that $(R,\langle\pi\rangle,\kappa)$ is a DVR such that $\kappa$ is quadratically closed of characteristic $\neq 2$, and set $f=\pi^{2}+x^{2}\in R[[x]]$. We show directly that $f$ is $2$-determined with respect to $\mathfrak{M}=\langle x,\pi\rangle$. Given some $g\in R[[x]]$ such that $f-g\in\mathfrak{M}^{3}$, we can write $g=x^{2}+\pi^{2}+h$ where $h\in\langle x,\pi\rangle^{3}$, and so $h=\alpha_{1}x^{3}+\alpha_{2}x^{2}\pi+\alpha_{3}x\pi^{2}+\alpha_{4}\pi^{3}$ for some $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}\in R[[\underline{x}]]$. Therefore we have that
@@
Observe that both $1+\alpha_{1}x+\alpha_{2}\pi$ and $1+\alpha_{3}x+\alpha_{4}\pi$ are units. Therefore, since $R$ is a DVR (and therefore Henselian) and since $\kappa$ is quadratically closed of characteristic $\neq 2$, there exists some unit $\beta\in R[[\underline{x}]]$ such that $\beta^{2}=(1+\alpha_{1}x+\alpha_{2}\pi)(1+\alpha_{3}x+\alpha_{4}\pi)^{-1}$. So, if we define the automorphism $\varphi$ by setting $\varphi(x)=\beta x$ then we can conclude that $\varphi(f)=(1+\varphi(\alpha_{3})\beta x+\varphi(\alpha_{4})\pi)g$ with $1+\varphi(\alpha_{3})\beta x+\varphi(\alpha_{4})\pi$ being a unit, and thus $f\overset{\mathcal{K}}{\sim}g$, as desired.
@@
Determinacy over local rings that are algebras over an algebraically closed field of characteristic zero have been studied by Belitskii and Kerner in [BK12, BK16b, BK19], and in [BGK22], Boix et al. presented a "characteristic-free" approach to determinacy over local ring by looking at group actions on modules over local rings. In addition, the study of the group $\mathcal{D}_{I}$ where $I$ is an ideal in the ring of power series over $\mathbb{C}$ was presented in [PEL88].
@@
  • The $R$-homomorphism $\varphi$ defined by $\varphi(x_{j})=x_{j}+b_{j}$ for every $j$ defines an $R$-automorphism of $R[[\underline{x}]]$.
@@
Note that we can rewrite Lemma 2.9 as $\operatorname{ord}_{I}(\partial_{i}(f))\geq\operatorname{ord}_{I}(f)-1$ for every ideal $I\subset R[[\underline{x}]]$. If $R$ is a ring of either mixed or positive characteristic then the order of the derivative can substantially change and is unbounded from above. For example, if $R$ is of mixed characteristic $(0,p)$ and $f(x)=x^{p^{r}}\in R[[x]]$ for some $r>1$ then we have that $\text{ord}(\partial_{x}(f))=\text{ord}(f)+r-1$.
@@
    (Implicit Function Theorem): Given a sequence of power series $f_{1},\dots,f_{m}\in R[[\underline{x},y_{1},\dots y_{m}]]$ such that $f_{i}(\underline{0})=0$ for every $i$ and
@@
    then we can find a sequence of power series $g_{1},\dots,g_{m}\in\mathfrak{M}\subset R[[\underline{x}]]$ for which we have that $f_{i}\left(x_{1},\dots,x_{n},g_{1},\dots g_{m}\right)=0$ for every $i$.
@@
    (Inverse Function Theorem): Let $f_{1},\dots,f_{n}\in\mathfrak{M}\subset R[[\underline{x}]]$. Then the Jacobian matrix $Df=\left[\frac{\partial f_{i}}{\partial x_{j}}\left(0\right)\right]_{i,j}\in\operatorname{Mat}_{n}\left(R\right)$ is an invertible matrix if and only if the unique $R$-homomorphism $\varphi\colon R[[\underline{x}]]\to R[[\underline{x}]]$ defined by $\varphi\left(x_{i}\right)=f_{i}$ defines an $R$-automorphism.
@@
    Since $R[[\underline{x}]]$ is a complete local ring, then by Lemma 10.153.9. in [STA18] (or by corollary 1.9 in [LW12]) it must be Henselian. Thus, given $f_{1},\dots,f_{m}\in R[[\underline{x},y_{1},\dots,y_{m}]]=R[[\underline{x}]][[y_{1},\dots,y_{m}]]$ such that satisfy the condition above, if we choose to define $\overline{f_{i}}=f_{i}\,\operatorname{mod}\,\mathfrak{M}\in\kappa[[y_{1},\dots,y_{m}]]$ for every $i$ we must have that $\overline{f_{i}}\in\langle y_{1},\dots,y_{m}\rangle\subset\kappa[[\underline{y}]]$ and that the Jacobian matrix of $(\overline{f_{1}},\dots,\overline{f_{m}})$ is an invertible matrix over $\kappa$. Therefore, by the Implicit function theorem over $\kappa$ and by the Multi-dimensional version of Hensel’s lemma (see, for example, Section 4 in [KUH00], Section 7 in [ZP78], or Section 1.2 of Chapter I in [GLS07]), we can find some unique $\overline{g_{1}},\dots,\overline{g_{m}}$ such that $\overline{f_{i}}\left(x_{1},\dots,x_{n},\overline{g_{1}},\dots,\overline{g_{m}}\right)=0$ for every $i$ that uniquely lift to some $g_{1},\dots,g_{m}\in R[[\underline{x},y_{1},\dots y_{m}]]$ such that $f_{i}\left(x_{1},\dots,x_{n},g_{1},\dots,g_{m}\right)=0$ for every $i$, as desired.
@@
In addition, if we have that $\varphi$ is an $R$-isomorphism then in fact we can choose the $R$-homomorphism $\Phi$ to be an $R$-isomorphism as well with $\Phi(I)=J$.
@@
Note that Lemma 2.13 and Lemma 2.14 give us analogues to some classical results in the field case. The first item of Lemma 2.13 is a generalized version of [TOU68] (for more details, see Section 3.1 in [HR13], Section 3 in [BK16a], or [KER24]) and the second item is a version of Corollary 3.3.7 in [DP13]. Lemma 2.14 is a generalized version of Lemma 1.14 and Lemma 1.23 in [GLS07] and of Lemma 1.3.5. in [BOU09].
@@
then we have that $F(\underline{x},\varphi_{i}(\underline{x}),u_{i})=f-u_{i}g\circ\varphi_{i}\to 0$. So by the Artin’s strong approximation theorem over local rings (see, for example, Theorem 3.3. in [RON15], which in turn cites [SCH80, SCH83]) we have that there exists some automorphism $\varphi$ and some unit $u$ such that $u_{i}\varphi_{i}(\underline{x})-u\varphi(\underline{x})\in\mathfrak{M}^{i}$ and $F(\underline{x},\varphi(\underline{x}),u)=0$, which gives us that $f=ug(\varphi)$. ∎
@@
Artin’s approximation theorem that we use in the proof of Lemma 2.16, which is based upon Artin’s original result in [ART69], is an essential result in the theory of isolated singularities over fields. For more information and details, see [BBK20, HR13].
@@
The proof of Lemma 2.18 is similar to that of Lemma 1.2.7. in [BOU09] and to that of Lemma 2.10. in [GLS07].
@@
which must be contained in $\mathfrak{M}^{N+2}$. Now, we can repeat this process inductively (using the notations of Lemma 2.7) to get collection of sequences $\{\{b_{i,j}\}_{i=1}^{\infty}\colon j=1,\dots,n\}$ such that $b_{i,0}\in\mathfrak{M}^{N+i-1}$ and $b_{i,j}\in\mathfrak{M}^{N+i}$ for every $j=1,\dots,n$. Therefore by setting $\varphi_{i}(x_{j})=x_{j}+b_{i,j}$ for every $i,j$, we get that $g-v_{i}\cdot\varphi_{i}(f)\in\mathfrak{M}^{N+1+i}$ and that $\varphi_{i}\in\mathcal{D}_{I}$, as desired. ∎